### v2

# Multi-Margin Loss: Proposal and Application in Recommender Systems

Abstract.
Recommender systems guide users through vast amounts of information by suggesting items based on their predicted preferences. Collaborative filtering-based deep learning techniques have regained popularity due to their simplicity, using only user-item interactions. Typically, these systems consist of three main components: an interaction module, a loss function, and a negative sampling strategy. Initially, researchers focused on enhancing performance by developing complex interaction modules with techniques like multi-layer perceptrons, transformers, or graph neural networks. However, there has been a recent shift toward refining loss functions and negative sampling strategies. This shift has increased interest in contrastive learning, which pulls similar pairs closer while pushing dissimilar ones apart. Contrastive learning involves key practices such as heavy data augmentation, large batch sizes, and hard-negative sampling, but these also bring challenges like high memory demands and under-utilization of some negative samples. The proposed Multi-Margin Loss (MML) addresses these challenges by introducing multiple margins and varying weights for negative samples. MML efficiently utilizes not only the hardest negatives but also other non-trivial negatives, offering a simpler yet effective loss function that outperforms more complex methods, especially when resources are limited. Experiments on two well-known datasets showed MML achieved up to a 20% performance improvement compared to a baseline contrastive loss function with fewer negative samples.
Recommender systems, Loss function
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Computing methodologies Neural networks
1. Introduction
Figure 1. Three major components for a deep-learning based CF model: (i) interaction module/encoder, (ii) loss function and (iii) negative sampling strategy. Given a representation space, user interacts with an item (e.g., purchases a product), shown with $p$, and from the remaining negative items $n$, four negatives, $n_{1}-n_{4}$, are sampled. Then, the sampled negative items, positive item and user representations are used by the interaction encoding and loss function modules for the computations.
[…]
Three major components for a deep-learning based CF model: (i) interaction module/encoder, (ii) loss function and (iii) negative sampling strategy.
Numerous web applications, such as YouTube, Amazon, Spotify, use recommendation systems to help users navigate the vast amounts of content (Park et al., 2023). These systems suggest items, such as videos, products, and music tracks, by estimating user preferences (Massa and Avesani, 2007; Tavakolifard and Almeroth, 2012). In the literature, recommender systems employ various approaches, including traditional methods like collaborative filtering, content-based filtering, matrix factorization(Pan et al., 2008; Ye et al., 2010; Li et al., 2015; He et al., 2016) as well as newer techniques such as deep learning(Ozsoy, 2016; Vasile et al., 2016; He et al., 2017; Musto et al., 2018), reinforcement learning (Zheng et al., 2018; Shi et al., 2021), language models-based methods (Sun et al., 2019; Fan et al., 2023b). Recently, collaborative filtering-based deep learning techniques regained popularity due to their simplicity, efficacy and reliance only on user-item interactions (Mao et al., 2021a; Park et al., 2023). These methods are composed of three major components: interaction module/encoder, loss function and negative sampling strategy (Mao et al., 2021a), The Figure 1 simplifies these steps: Given a representation space, user interacts with an item (e.g., purchases a product), shown with $p$, and from the remaining negative items $n$, four negatives, $n_{1}-n_{4}$, are sampled. Then, the sampled negative items, positive item and user representations are used by the interaction encoding and loss function modules for the computations and making recommendations.
For years, researchers focused on improving interaction modules using various techniques such as multi-layer perceptrons (Covington et al., 2016; He et al., 2017), auto-encoders (Liang et al., 2018), attention networks (Chen et al., 2020a), transformers (Sun et al., 2019), graph neural networks (He et al., 2020; Ozsoy et al., 2020). Even though these models showed better performance results, they became more and more complex. On the other hand, in industrial recommender systems, highly efficient models are preferred because of their practical applicability. Recently, SimpleX (Mao et al., 2021a) showed that a simple interaction encoding model can outperform complex techniques with a well-designed loss function. This success has shifted focus in the recommender systems to loss functions and negative sampling strategies. Several recent studies analyze the impact of loss functions on recommendation performance (Wu et al., 2022; Li et al., 2023; Park et al., 2023), while others (Zhang et al., 2022; Wang et al., 2022; Lee et al., 2023; Sun et al., 2023; He et al., 2023; Wu et al., 2023; Zhang et al., 2024) develop new loss functions to address challenges like popularity bias and robustness. These efforts have increased interest in contrastive learning in the recommender systems domain.
The models using contrastive learning aim to learn representations by pulling similar pairs closer while pushing dissimilar pairs apart. The learning process is highly relevant to the designed contrastive loss functions (Wang and Liu, 2021). For an efficient contrastive learning system, three key components have been identified (Weng, 2021): heavy data augmentation, large batch size and hard-negative sampling. While using larger batch sizes and more negative samples introduce more diverse and ‘hard’-negative samples to the system, they require more memory as a side-effect. A common approach is to use a margin parameter to distinguish ‘hard’-negatives from the rest. However, this approach sets only a single threshold, failing to fully utilize all available negative samples efficiently.
In this work, we hypothesise that using not only hardest negative instances but all non-trivial ones allow model to generalize more efficiently, even with smaller number of negative samples or batch sizes. Furthermore, assigning different weights to varying levels of ”hardness” in negative instances—such as giving higher weight to the hardest negatives and relatively lower weights to easier ones—can enhance performance. For this purpose, we propose Multi-Margin Loss (MML), which is a simple yet efficient loss function, that uses multiple margins instead of a single margin and assigns different weights to different levels of negative instances.
The main contributions of this work are summarized as follows:
  • We propose the Multi-Margin Loss (MML) function, which introduces multiple margins and weights for negative instances. This structure allows us to consider not only the hardest negatives but also other non-trivial ones, like ‘semi-hard’ and ‘semi-easy’ negatives. The ability to assign different weights enables tuning the importance assigned to each level of negative samples.
  • We conducted experiments on two well-known and frequently used recommender system datasets and compared MML’s performance to state-of-the-art loss functions.
  • Our results show that (i) with a large negative sample size (e.g., 800), MML performs as well as state-of-the-art contrastive loss functions, and (ii) with smaller sample sizes (e.g., 10 or 100), MML achieves up to a 20% performance improvement over baseline contrastive loss functions.
The rest of the paper is structured as follows: Section 2 provides background on recommender systems, contrastive learning and loss functions. Section 3 introduces the Multi-Margin Loss (MML). Section 4 covers the experimental setup and evaluation results. Finally, Section 5 discusses the conclusion and future directions.
2. Related Work
This section provides background on recommender systems and collaborative filtering, an overview of loss functions used in these systems, and more details on contrastive loss functions.
Figure 2. (a) Margin $m$ defines a radius, which makes the model to pay attention to harder-negatives. In Contrastive Loss, Eq.1, $m1$ is used, where radius is defined relative to user representation. In Triplet Loss, Eq. 3, $m2$ is used, where radius is defined relative to positive item. (b) Proposed Multi-Margin Loss (MML), Eq. 7, filters not only hardest-negatives but also other non-trivial negatives and assigns different weights for each level. It uses multiple margins, $m1,m2,m3$, and multiple weights,$w1,w2,w3$.
[…]
Margin $m$ defines a radius, which makes the model to pay attention to harder-negatives.
2.1. Recommender Systems and Collaborative Filtering-based models
Recommender systems (RS) aim to estimate users’ preferences and recommend items based on historical user-item interaction data. RS can use various approaches, such as traditional collaborative filtering, content-based filtering, matrix factorization (Pan et al., 2008; Ye et al., 2010; Li et al., 2015; He et al., 2016), deep learning (Ozsoy, 2016; Vasile et al., 2016; He et al., 2017; Musto et al., 2018), reinforcement learning (Zheng et al., 2018; Shi et al., 2021), and language model-based methods (Sun et al., 2019; Fan et al., 2023b). Recently, collaborative filtering-based deep learning techniques have regained popularity due to their simplicity, efficacy, and reliance only on user-item interactions (Mao et al., 2021a; Park et al., 2023). Collaborative Filtering (CF) models use only interaction history without any other contextual information and consist of three major components (Mao et al., 2021a; Park et al., 2023), as shown in Figure 1.
  • Interaction encoding: Interaction encoder aims to learn embedding of each user and item. It has been primary focus in the field for many years. To improve it, various techniques such as vector embeddings (Ozsoy, 2016; Vasile et al., 2016), multi-layer perceptrons (Covington et al., 2016; He et al., 2017), auto-encoders (Liang et al., 2018), attention networks (Chen et al., 2020a), transformers (Sun et al., 2019), graph neural networks (He et al., 2020; Ozsoy et al., 2020), reinforcement learning techniques (Zheng et al., 2018; Shi et al., 2021) have been employed. Despite better performance, these methods became increasingly complex. In industry, however, efficient models are in higher demand due to their practical application (Mao et al., 2021a).
  • Loss function: The commonly used loss functions in RS were Binary Cross-Entropy (BCE) loss (He et al., 2017), Bayesian Personalized Ranking (BPR) loss (Rendle et al., 2009) or mean square error (MSE) loss (Chen et al., 2020b). In the other research domains, such as computer vision, natural language processing, many other loss functions have been developed and applied, which utilise both positive and negative instances and contrastive learning approaches. Inspiring from their success, in the last few years more focus is given to the loss functions in recommender applications.
  • Negative sampling strategy: Negative sampling aims sample subset of negative instances from a larger set (Yang et al., 2024). In RS, there are vast amount of negative or unobserved instances compared to positive ones. In order to improve training efficiency negative sampling techniques are applied. There are various negative sampling strategies, from random negative sampling (RNS) (Pan et al., 2008; Ding et al., 2019), popularity-based negative sampling (PNS) (Barkan and Koenigstein, 2016; Yu et al., 2017; Lian et al., 2020) to hard negative sampling (hard NS) (Zhang et al., 2013; Rendle and Freudenthaler, 2014), and GAN-based negative sampling (GAN-based NS)(Jin et al., 2020).
In the last few years many researchers in recommender systems focused more on loss functions. In Sections 2.2 and 2.3, more details on loss functions used in recommender systems, specifically contrastive loss functions are presented.
2.2. Overview of Loss Functions used in Recommender Systems
In general, there are three common types of loss functions in recommender systems:
  • Point-wise loss function measures the distance between the predicted and the expected labels. Each instance is used independent from the other instances, i.e., either a positive or a negative instance is used at a time. Example point-wise loss functions used in RS are Binary Cross Entropy (BCE) (He et al., 2017), Mixed-centric (MCL) (Gao et al., 2022) and User Interest Boundary (UIB) (Zhuo et al., 2022) losses.
  • Pair-wise loss function checks a pair of items at a time and aims to decide the optimal ordering of the pair and compare it to the ground truth. Example pair-wise loss functions used in RS are Bayesian Personalized Ranking (BPR) loss (Rendle et al., 2009), Collaborative Metric Learning (CML) (Hsieh et al., 2017) and Symmetric Metric Learning (SML) (Li et al., 2020).
  • Set-wise loss functions are generalization of the pairwise loss functions (Park et al., 2023). Instead of a pair of positive and negative items, it considers multiple negative items for a single positive item. Example set-wise loss functions used in RS are Cosine Contrastive Loss (CCL) (Mao et al., 2021a), Sampled Softmax (SSM) (Wu et al., 2022), Bias-aware Contrastive (BC) (Zhang et al., 2022) losses.
Another direction in recommender systems is the use of automated machine learning (AutoML) techniques to explore loss functions (Chen et al., 2024). SFL (Liu and Lai, 2020) and AutoLoss (Zhao et al., 2021) aim to combine loss functions automatically, while AutoLossGen (Li et al., 2022) generates new loss functions from basic mathematical operators. We believe that creating well-designed loss functions and applying AutoML techniques for their generation and exploration will complement each other, enhancing the overall performance of recommender applications.
Recently, SimpleX (Mao et al., 2021a) demonstrated that a well-designed loss function can enable a simple model to outperform more complex techniques, sparking interest in designing loss functions for recommender systems. Some studies have analyzed the impact of loss functions on recommendation performance (Wu et al., 2022; Li et al., 2023; Park et al., 2023; Wu et al., 2023), while others (Zhang et al., 2022; Wang et al., 2022; Lee et al., 2023; Sun et al., 2023; He et al., 2023; Zhang et al., 2024) have developed new loss functions to address challenges like popularity bias. These investigations have highlighted contrastive learning as an increasingly popular trend.
2.3. Contrastive Learning
In the machine-learning literature, especially for unsupervised methods, contrastive learning drew increased attention due to their better performances (Wang and Liu, 2021). The models using contrastive learning aim to learn representations by making similar pairs closer to each other while dissimilar one are separated. The learning process is highly relevant to the designed contrastive loss functions (Wang and Liu, 2021).
Scientific Notation
Given a set of users, $U$, and set of items, $I$, user-item interaction can be represented as $R\in\{0,1\}^{|U|x|I|}$. In this setting, $r_{ui}=1$ and $r_{ui}=0$ indicate positive and negative interactions, respectively. Note that negative interaction can have two meanings: (1) The item was shown to the user and the user did not prefer (not clicked/purchased/…) that item or (2) The item was not shown to the user at all. The items in positive interactions are named as positive instances: $I_{u}^{+}=\{i\in I|r_{ui}=1\}$ and, similarly, items in negative interactions are named as negative instances: $I_{u}^{-}=\{i\in I|r_{ui}=0\}$.
In Collaborative Filtering, only user-item interactions are used as the input features. Various loss functions, indicated as $f$, are used in the literature for training. For the computations, they use predicted relevance $s(u,i)$, predicted distance $d(u,i)$ and actual label $l(u,i)$ of the user-item interactions. Some of the loss functions use margin $m$ and weight $w$ parameters.
Contrastive learning
: In this section, commonly used contrastive loss functions are summarized:
  • Contrastive loss (Chopra et al., 2005): Contrastive loss is one of the earliest loss function defined for contrastive learning. Having pair of user $u$ and item $i$ representation (i.e., $i\in I+$ or $i\in I-$), the goal is to make representation of positive items closer and representation negative items further away from the user representation. The equation for contrastive loss can be written as follows:
    (1) $\begin{split}L_{Contrastive}=\dfrac{1}{2}\sum_{{x}\in{I}}Y_{i}d(u,i)^{2}{\ +}\\ (1-Y_{i})max(0,m-d(u,i))^{2}\end{split}$
    In the equation, $Y_{i}$ indicates function which returns 1 if item is positive (i.e., $i\in I+$) and 0 otherwise. Margin $m$ defines a radius, which makes the model to pay attention to hard-negatives, which are negative instances which has the lowest distance (i.e., most similar) to the user representation. The idea is also presented in Figure 2a. In the figure, margin $m1$ is used for filtering negative samples. As the result, only $n1$, which is in the hard-negatives area, and positive item $p$ are used for the computations.
    One of the first application of contrastive loss in recommender systems belongs to SimpleX (Mao et al., 2021a). In their work, they have proposed Cosine Contrastive Loss (CCL), which utilizes a similar idea, but with a set of negative items and using cosine similarity instead of distance. The Eq.2 for CCL (Mao et al., 2021a) can be written as follows:
    (2) $\begin{split}L_{CCL}=\sum_{{i}\in{I+},{N}\subset{I-}}1-s(u,i){\ +}\\ \dfrac{w}{N}\sum_{{j}\in{N}}max(0,s(u,j)-m)\end{split}$
  • Triplet loss (Schroff et al., 2015): Triplet loss is another commonly used contrastive learning loss function. It was proposed in the FaceNet (Schroff et al., 2015) for a face recognition task. Having triplets of user $u$, positive item $i\in I+$ and negative item $j\in I-$, triplet loss aims to minimize distance between user and positive item and maximize the distance between user and negative item, at the same time (Weng, 2021). The equation for triplet loss can be written as follows:
    (3) $L_{Triplet}=\sum_{i\in I+,j\in I-}max(0,d(u,i)^{2}-d(u,j)^{2}+m)$
    In the equation, margin $m$ defines a radius relevant to distance of the positive item. As the result: (i) If the distance of negative item to the user is less than the distance of the positive item, i.e., $d(u,n)<d(u,p)$, then those items are the hardest ones and included in the computations. (ii) If the distance of negative item is between the distance of the positive item and summation of distance of positive item and margin, i.e., $d(u,p)<d(u,n)<d(u,p)+m$, then those items are the semi-hard negatives and also included in the computations. The idea is also presented in Figure 2a. In the figure, margin $m2$ is used for filtering negative samples. As the result, $n1$ and $n2$, which are in the radius of positive item $p$, are used for the computations.
  • InfoNCE loss (Oord et al., 2018): Noise Contrastive Estimation (NCE) (Gutmann and Hyvärinen, 2010) aims to distinguish the target data from noise. The InfoNCE loss uses this idea to rank the positive samples higher from a set of positive and negative samples. The equation for InfoNCE loss can be written as follows:
    (4) $L_{InfoNCE}=-\sum_{i\in I+}log\dfrac{exp(s(u,i))}{exp(s(u,i))+\sum_{j\in N}{exp(s(u,j))}}$
    AdvInfoNCE(Zhang et al., 2024) further improves InfoNCE loss by adaptively exploring and assigning a difficulty level to each negative instance in an adversarial manner.
  • Bilateral SoftMax Loss (Wu et al., 2023): Bilateral SoftMax Loss (BSL) (Wu et al., 2023) is one of the latest loss functions proposed in recommender systems. It extends Softmax loss (SL) (Wu et al., 2022), which normalizes model predictions and optimizes the probability of positive instances versus negative ones. BSL applies the same techniques as in SL on both positive and negative instances. As the result, the model becomes more robust to noisy data. Its equation can be written as follows:
    (5) $\begin{split}L_{BSL}=-\tau_{1}log\mathbb{E}_{i\sim P+}\dfrac{exp(s(u,i))}{\tau_{1}}{\ +}\\ \tau_{2}log\mathbb{E}_{j\sim P-}\dfrac{exp(s(u,j))}{\tau_{2}}\end{split}$
    In the equation, $P$ indicates the distribution of instances, $\mathbb{E}$ indicates expectation, $\tau$ values indicate temperatures of components.
3. Proposed Loss Function
In this section, we introduce the Multi-Margin Loss (MML). We start by discussing the generalization of contrastive loss functions, then we explain the rationale behind the MML function. Finally, we provide a detailed description of the MML function itself.
3.1. Generalized Contrastive Loss
From the loss functions summarized in related work section, we can generalize the contrastive loss as follows:
(6) $\begin{split}L_{\begin{subarray}{c}Generalized\\ Contrastive\end{subarray}}=\sum_{i\in I+,N\subset I-}w_{p}f(u,i){\ +}\\ \dfrac{w_{n}}{N}\sum_{j\in N}max(0,f(u,i,j,m))\end{split}$
In the equation,
  • A set-wise loss is used for generalization. Set-wise loss computes the loss over a positive instance and a set of N negative samples. If only a single negative item is used, it reduces to pair-wise loss.
  • Distinct weights for positive and negative instances are used. Previously, weight for the positive instance, $w_{p}$, is set either to 1.0 as in Base Contrastive Loss, Eq.1 or in CCL, Eq.2, or to 0.0 as in Triplet Loss, Eq.3.
  • A generic function $f(u,i,j,m)$ used for representing various combinations of margin $m$ and positive and negative instances, $i\in I^{+}$ and $j\in I^{-}$, respectively.
$f(u,i,j,m)$ can be any function defined previously: For example, InfoNCE, Eq.4, do not use margin at all; $f(u,i,j,m)=s(u,j)$. While CCL, Eq.2, uses similarity between the user and the negative instance and margin; $f(u,i,j,m)=s(u,j)-m$, Base Contrastive Loss, Eq.1, uses distance function with margin; $f(u,i,j,m)=m-d(u,j)$. Triplet Loss, Eq.3, uses both positive and negative items with margin; $f(u,i,j,m)=d(u,i)-d(u,j)+m$, where margin is relative to the representation of the positive item; i.e., distance from the positive item, $d(u,i)-d(u,j)$ is used.
3.2. Motivation for Multi-Margin Loss
For an efficient contrastive learning system, Weng et al.(Weng, 2021) identify three key components: (i) Heavy Data Augmentation: Data augmentation techniques create noisy versions of positive instances and are critical for learning generalizable embedding features. (ii) Large Batch Size: In order to cover a diverse and challenging negative samples, it is necessary to have larger batch sizes. (iii) Hard-Negative Mining: Hard negatives are the instances which are similar to the anchor item, i.e., user for recommender systems, but which are negative, i.e., user did not like that item. One solution for incorporating hard negatives is to use large batch sizes or more negative samples, which implicitly introduce more hard negative samples to the system. However, this approach suffers from higher memory usage as a side effect. A more efficient and commonly used solution is to use a margin $m$ for the computations. With a margin $m$, it is possible to distinguish hard negatives from the rest.
When examining the key components of contrastive learning, we recognized the importance of effectively utilizing negative samples while prioritizing the more challenging ones. Instead of relying only on larger batch sizes to introduce more hard negatives, which can be memory-intensive, we propose that using not just the most challenging instances but all non-trivial examples can lead to better model generalization, even with fewer negative samples or smaller batch sizes. We propose assigning different weights based on the ”hardness” of negative samples, with higher weights given to the hardest negatives and relatively lower weights to easier ones. This prioritization helps the model to distinguish between different levels of ”hardness” among negative samples, effectively using a smaller number of negative samples. Furthermore, considering all non-trivial negative examples enhances generalization, ensuring the model learns from a diverse set of negative samples.
Training a recommender model using contrastive learning helps distinguish between similar and dissimilar pairs and learn representations that capture meaningful relationships between users and items. Better representation of users and items can improve recommendation performance by generalizing well even with limited interaction data, understanding a broader spectrum of item similarities and dissimilarities, and ensuring that the negative pairs represent diverse and unbiased samples. To address these considerations, we propose the Multi-Margin Loss, which is detailed in Section 3.3.
3.3. Multi-Margin Loss
The proposed Multi-Margin Loss (MML) function introduces multiple margins and weights for negative instances. Having various margins allows us to consider not only hardest-negatives but also other non-trivial negatives; i.e., ’semi-hard’-negatives, ’semi-easy’-negatives. Utilization of weights allows us to tune the importance for each level of negatives.
In this work, we used a similar approach to CCL (Mao et al., 2021a) and extended the generalized contrastive loss. Even though it is possible to extend any loss functions using multi-margins, for the experiments CCL loss (Mao et al., 2021a) is used as the baseline and we extended the generalized contrastive loss. Overall MML equation can be written as follows:
(7) $\begin{split}L_{MML}=\sum_{i\in I+,N\subset I-}w_{p}(1-s(u,i)){\ +}\\ \dfrac{1}{N}\sum_{j\in N}\sum_{k\in|M|}w_{k}max(0,s(u,j)-m_{k})\end{split}$
In the equation $|M|$ is the size of the margins defined. Each margin in $M$; $m_{k}\in M$, is assigned with a different weight,$w_{k}$. Similar to CCL (Mao et al., 2021a), (i) similarity is used for the computations: $f(u,i)=1-s(u,i)$ and $f(u,i,j,m_{k})=s(u,j)-m_{k}$ and (ii) margins are set relative to the representation of the user. The idea of MML is also presented in Figure 2b. MML filters not only hardest-negatives but also other non-trivial negatives and assigns different weights for each level. In the figure, it uses multiple margins, $m1,m2,m3$ $m1$, and multiple weights $w1,w2,w3$. As the result, not only hardest-negatives, e.g., $n1$, but other non-easy/non-trivial negatives, e.g.,$n2$ and $n3$, are selected and used in the computations.
Table 1. The statistics of the datasets
Yelp Gowalla
#Users 31,668 29,858
#Items 38,048 40,981
#Interactions 1,561,406 1,027,370
#Train 1,237,259 810,128
#Test 324,147 217,242
Density 0.00130 0.00084
Table 2. Performance comparison of MML and CCL on SimpleX and MF architectures with and without early-stopping (Number of negative samples = 100). Among the all settings, MML consistently delivers the best results on the Yelp and Gowalla datasets.
Model Loss Early-stop? Yelp18 Gowalla
Recall@20 NDCG@20 Recall@20 NDCG@20
SimpleX CCL (2023) Y 0.0634 0.0517 0.1622 0.1283
CCL (2023) N 0.0679 0.0555 0.1663 0.1319
MML (2024) - Proposed Y 0.0662 0.0539 0.1723 0.1358
MML (2024) - Proposed N 0.0706 0.0578 0.1787 0.1412
MF CCL (2023) Y 0.0601 0.0490 0.1632 0.1279
CCL (2023) N 0.0673 0.0551 0.1783 0.1426
MML (2024) - Proposed Y 0.0674 0.0550 0.1756 0.1370
MML (2024) - Proposed N 0.0702 0.0580 0.1816 0.1435
Table 3. Training-Runtime Comparison on MF model (Without early-stopping, 100 epochs). MML increases training time per epoch by at most 2 seconds on average.
#Negative samples Loss Yelp18 Gowalla
Time per epoch Ratio Total time Time per epoch Ratio Total time
800 CCL (2021) 77.5 secs - $\sim$129 mins 66 secs - 110 mins
MML (2024) - Proposed 78.5 secs +1.3% $\sim$130 mins 64 secs -3.0% $\sim$107 mins
100 CCL (2021) 35 secs - $\sim$58 mins 37 secs - $\sim$62 mins
MML (2024) - Proposed 35 secs 0.0% $\sim$58 mins 37.5 secs +1.4% $\sim$63 mins
10 CCL (2021) 31 secs - $\sim$51 mins 32 secs - $\sim$53 mins
MML (2024) - Proposed 33 secs +6.5% 55 mins 33.5 secs +4.7% $\sim$56 mins
4. Evaluation Results
This section presents the evaluation setup and the results.
4.1. Evaluation Setup
SimpleX (Mao et al., 2021a) demonstrated that even with a simple neural-network architecture, a well-designed loss function can achieve good performance results. For comparisons, SimpleX (Mao et al., 2021a), which proposes the CCL loss function, is used as the main benchmark. In addition to the CCL loss, five more loss functions are used for comparisons: (i) four classical loss functions: BPR loss, Pairwise Hinge loss, Softmax Cross-Entropy loss, and Mean Square Error loss; (ii) two recently proposed contrastive loss functions: CCL (Mao et al., 2021a) and BSL (Wu et al., 2023).
Datasets and Evaluation Metrics
Two datasets, which are commonly used in recommender systems, especially by recent neural-network-based CF models (Wang et al., 2019; Chen et al., 2020a; He et al., 2020; Mao et al., 2021a, b; Fan et al., 2023a; Zhou et al., 2023; Wang et al., 2024) are used: (i) Yelp2018 (Yelp, 2018) [^1] Yelp2018: https://huggingface.co/datasets/reczoo/Yelp18_m1 and (ii) Gowalla (Cho et al., 2011) [^2] Gowalla: https://huggingface.co/datasets/reczoo/Gowalla_m1. These datasets have already been split, pre-processed, and made publicly available. The statistics of the datasets are summarized in Table 1. Same as the SimpleX (Mao et al., 2021a), we used the datasets as is, without executing any preprocessing and/or filtering, such as filtering by popularity or cold-start. We consider all items not interacted with by a user as candidates and report the average results across all users.
In order to evaluate the performance of recommendations, we use commonly accepted metrics like Recall and NDCG. In this work, we specifically present the Recall@20 and NDCG@20 metrics for comparing performance, which are also used by our comparison partners, CCL (Mao et al., 2021a) and BSL (Wu et al., 2023).
(a) Yelp dataset
(b) Gowalla dataset
Figure 3. Performance of MF model under recent loss functions while using various number of negative-samples. Increasing number of negative-samples improves the performance for all loss functions. The proposed loss function, MML, outperforms existing baseline loss functions when the number of negative samples is 100 or fewer, suggesting it is more efficient in resource-constrained environments.
Implementation Details
As explained previously, the main comparison partners are SimpleX (Mao et al., 2021a), which proposes CCL, and BSL (Wu et al., 2023). For the computations P100 GPU is used, with a limited access quota of 30 hours per week. Consequently, the experiments primarily focused on enhancing performance with limited computational capacity.
For the implementation of MML, the code provided by SimpleX (Mao et al., 2021a) is expanded and shared online [^3] MML (Proposed) code: TO-BE-SHARED. The experiments involving CCL utilized the code from the SimpleX (Mao et al., 2021a) [^4] SimpleX (CCL) (Mao et al., 2021a) code: https://reczoo.github.io/SimpleX. For experiments related to BSL, we used the code provided in the BSL (Wu et al., 2023) paper [^5] BSL (Wu et al., 2023) code: https://github.com/junkangwu/BSL. Although the original code implementations are utilized, all the experiments have been re-run and the results are presented in the following sections. In the experiments, we used the same parameters as the SimpleX (Mao et al., 2021a) paper. For both the Yelp and Gowalla datasets, we used the Adam optimizer with $L_{2}$ regularization. The learning rate for both datasets was set to 1e-4, with an embedding dimension of 64 and 100 epochs. The batch size for the Yelp dataset was 512, with a regularizer of 1e-09, while for Gowalla, the batch size was 256, with a regularizer of 1e-08. We explored various numbers of negative samples, but results are presented using 100 negatives, unless otherwise specified.
The MML-specific parameters were determined through grid search across the parameter space. We explored $w_{p}$:$w_{n}$ ranging from [1:150] to [1:450] in increments of 50 for Yelp and from [1:500] to [1:1000] in increments of 100 for Gowalla. Initial margin values started at [0.9] and were expanded until [0.5, 0.6, 0.7, 0.8, 0.9] for both datasets. Margin weights were adjusted with priority given to the highest margin and reductions made proportionally. As a result, for Yelp, margins were set to [0.6, 0.7, 0.8, 0.9] with corresponding weights of [0.05, 0.15, 0.2, 0.6], and $w_{p}$:$w_{n}$ was set to [1:400]. For Gowalla, margins were set to [0.7, 0.8, 0.9] with weights of [0.10, 0.15, 0.75], and $w_{p}$:$w_{n}$ was set to [1:700]. We want to use AutoML tools for searching the right parameters as a future direction. For the CCL experiments, we used the original code and configuration, with the margin kept at 0.9. Unlike the original setup, we did not use early-stopping, as it was shown to produce better results for our datasets. Additionally, we did not apply learning-rate reduction to align with the BSL implementation, which doesn’t support this feature. In the BSL experiments, we retained the original code and configuration. The temperatures $\tau_{1}and\tau_{2}$ set as 0.11 and 1.10 for the Yelp dataset, and 0.08 and 1.22 for the Gowalla dataset. The original code specified 1000 epochs, but for consistency with the other experiments, we set the number of epochs to 100.
Table 4. Performance of MF model with different loss functions (Number of negative samples = 100). MML performs equally well in recall and NDCG on the Yelp dataset and achieves the best results in Recall on the Gowalla dataset.
Loss Yelp18 Gowalla
Recall@20 NDCG@20 Recall@20 NDCG@20
BPR Loss 0.0272 0.0216 0.0708 0.0509
Pairwise Hinge Loss 0.0340 0.0269 0.0758 0.0541
Softmax Cross-Entropy 0.0309 0.0246 0.0734 0.0543
Mean Square Error 0.0609 0.0502 0.1391 0.1171
CCL (2021) 0.0673 0.0551 0.1783 0.1426
BSL (2023) 0.0703 0.0580 0.1804 0.1521
MML (2024) - Proposed 0.0702 0.0580 0.1816 0.1435
4.2. Evaluation Results
In this section, we examine the performance of the proposed MML model in comparison to baseline loss functions on various architectures and parameter configurations.
Impact of model architecture and early-stopping
SimpleX (Mao et al., 2021a) showed that their proposed CCL loss function combined with relatively simple neural network architectures like SimpleX and Matrix Factorization (MF) outperforms other more complex models.
In the SimpleX implementation (Mao et al., 2021a), the authors utilized an early-stopping mechanism, which stops training when the performance of the model on the validation set no longer improves. Table 2 presents a performance comparison of the CCL and MML loss functions on the SimpleX and MF architectures, with and without the use of early-stopping. With early-stopping, the SimpleX architecture concluded after 33 epochs with CCL and 19 epochs with MML, while the MF architecture ended after 19 epochs for both loss functions.
The table demonstrates that in our experiment omitting early-stopping results in better performance for both the CCL and MML loss functions, as well as for both the SimpleX and MF architectures. Among the all settings, MML consistently delivers the best results on the Yelp and Gowalla datasets. For the Yelp dataset, without early-stopping, MML outperforms CCL by 3.98% in recall and 4.14% in NDCG on SimpleX and by 4.31% in recall and 5.26% in NDCG on MF. On the Gowalla dataset, the gains with MML are 7.46% in recall and 7.05% in NDCG on SimpleX, while on MF, the improvement is 1.85% in recall and 0.63% in NDCG. We attribute the relatively lower performance gain on Gowalla dataset to the sparsity of the Gowalla dataset. In this dataset, MML can effectively learn from the available negative instances by considering all non-trivial negatives. However, due to the sparsity of the dataset, MML may struggle to learn sufficiently from the positive instances.
Training Runtime Comparison
We compare the training runtime of the CCL and MML loss functions using MF architecture in Table 3. The table shows the average time spent per epoch, change ratio and total training time on the Yelp and Gowalla datasets using varying numbers of negative samples. All the experiments use the same setup described in the Implementation Details in Section 4.1.
The Table 3 shows that for both CCL and MML loss functions the time spent per epoch increases as the number of negative samples increases. Training a model with 800 negative samples takes about 2 hours, but it takes around 1 hour with 100 negative samples. Comparing CCL and MML shows that using multiple margins has a relatively small impact, with the training time per epoch increasing by no more than 2 seconds on average. This small increase in training time, while using negative samples more efficiently, makes MML a promising loss function for recommender systems.
Impact of number of negative samples
One of the main objective of MML is to utilize better all the negative samples, while keeping batch size and number of negatives limited; hence reducing the memory requirements during training. To evaluate this, we compared the performance of recent loss functions against MML across various numbers of negative samples, illustrated in Figure 3.
The figure shows that MML outperforms CCL when there’s a large number of negative samples, though it’s not as effective as BSL. However, when the number of negative samples is 100 or fewer, MML consistently beats the baseline loss functions. When number of negative instances is set to 10, MML outperforms baseline CCL by 19.5% in recall and 20.13% in NDCG on Yelp dataset and by 12.99% in recall and 15.07% in NDCG on Gowalla dataset. An interesting observation from the figure is that on the Gowalla dataset, MML outperforms both CCL and BSL in terms of recall but falls behind BSL on the NDCG metric. This discrepancy can be attributed to MML’s effectiveness in learning from all available non-trivial negative instances, while the sparsity of the dataset makes it challenging for MML to gain sufficient insights from positive instances. Since all compared loss functions achieve relative stability when the number of negatives is 100, we used this value throughout the experiments.
Performance comparison of MML against different loss functions
The proposed MML loss function is compared with various loss functions using the MF architecture as the base model. The Table 4 lists the comparison results of MML with other loss functions. In the table, the first group contains the traditional loss functions commonly used in recommender systems. The second group contains the loss functions proposed recently and which are the state-of-the-art.
The table reveals that BSL is a strong baseline, outperforming MML in terms of recall on the Yelp dataset and in NDCG on the Gowalla dataset. Conversely, MML performs equally well in recall and NDCG on the Yelp dataset and achieves the best results in recall on the Gowalla dataset. The best performing traditional loss function was MSE for the configs we used. MML outperforms this loss function by 15.27% in recall and 15.54% in NDCG on Yelp dataset and 30.55% in recall and 22.54% in NDCG on Gowalla dataset.
5. Conclusion
Recommender systems guide users by estimating their preferences through methods like collaborative filtering and deep learning. Although sophisticated interaction modules using complex neural networks are common, simpler models with effective loss functions can sometimes achieve better or comparable performance. Recently, there has been a shift in focus toward loss functions and negative sampling strategies. This shift has led to a greater emphasis on contrastive learning, which brings similar pairs closer while separating dissimilar ones.
In order to address the challenges of efficient contrastive learning, a new loss function called Multi-Margin Loss (MML) is proposed, which uses multiple margins to manage different levels of ‘hardness’ among negative samples and introduces weights to assign varying importance to them. The experimental results executed on two real-world datasets showed that MML performs as well as or better than other loss functions depending on the size of negative sample sets. The results demonstrated that MML achieved up to a 20% performance improvement compared to a baseline contrastive loss function when fewer number of negative samples are used.
While MML has shown significant performance improvement, especially with smaller sets of negative samples, it can be further enhanced by incorporating ideas from other loss functions that address challenges in recommender systems, such as popularity bias, diversification and robustness. In addition to recommender systems, contrastive learning has proven to be effective in other domains. Future work could explore applying MML and other loss functions from recommender systems to these other domains.